# Possible Liquid-Nitrogen-Temperature Superconductivity Driven by Perpendicular Electric Field in the Single-Bilayer Film of La3Ni2O7 at Ambient Pressure

### v2

@@ III.2 The two-orbital model
where $\delta_{\mu\alpha}=\left\langle{}b_{i\mu\alpha}b^{{\dagger}}_{j\mu\alpha}\right\rangle$ since holon condense at zero temperature. Under the electric field, we have $\delta_{bz}=0$ and $\delta_{tz}$, $\delta_{tx}$ and $\delta_{bx}$ are solved in a self-consistent manner by adjustment to onsite energies $\epsilon_{\alpha}$ (See SI [146] for more details). The mean-field order parameters are represented by
@@
$\Delta^{\bm{x}}_{\mu x}=\frac{1}{2N}\sum_{\langle i,j\rangle}\Delta_{ij,\mu x},\ \Delta^{\perp}_{\alpha}=\frac{1}{N}\sum_{i}\Delta^{\perp}_{i\alpha}.$
Notably, the spin-exchange $\tilde{J}_{\perp}$ of the Hamiltonian in Eq. (2) doesn’t produce a hopping term $\chi_{x}^{\perp}$ in Eq. (5), which is the feature of such a bilayer system. Without interlayer hopping, a small interlayer spin-exchange $J_{\perp}$ leads to $\langle\chi^{\perp}\rangle\approx 0$.
@@
$\langle\chi_{\mu x}\rangle=\frac{1}{N}\sum_{k}\epsilon(\bm{k})\left(\langle f^{\dagger}_{k\mu x\uparrow}f_{k\mu x\uparrow}\rangle+\langle f^{\dagger}_{-k\mu x\downarrow}f_{-k\mu x\downarrow}\rangle\right),$
@@
$\langle\Delta^{\bm{x}}_{\mu x}\rangle^{*}=\frac{1}{N}\sum_{k}2\cos{(k_{x})}\langle f^{\dagger}_{k\mu x\uparrow}f^{\dagger}_{-k\mu x\downarrow}\rangle,$

### v3

@@ III.2 The two-orbital model
where $\delta_{\mu\alpha}=\left\langle{}b_{i\mu\alpha}b^{{\dagger}}_{j\mu\alpha}\right\rangle$ since holon condense at zero temperature. Under the electric field, we have $\delta_{bz}=0$ and $\delta_{tz}$, $\delta_{tx}$ and $\delta_{bx}$ are solved in a self-consistent manner by adjustment to onsite energies $\epsilon_{\alpha}$ (See SI [SM] for more details). The mean-field order parameters are represented by
@@
$\Delta^{\mathbf{x}}_{\mu x}=\frac{1}{2N}\sum_{\langle i,j\rangle}\Delta_{ij,\mu x},\ \Delta^{\perp}_{\alpha}=\frac{1}{N}\sum_{i}\Delta^{\perp}_{i\alpha}.$
Notably, the spin-exchange $\tilde{J}_{\perp}$ of the Hamiltonian in Eq. (2) doesn’t produce a hopping term $\chi_{x}^{\perp}$ in Eq. (S22), which is the feature of such a bilayer system. Without interlayer hopping, a small interlayer spin-exchange $J_{\perp}$ leads to $\langle\chi^{\perp}\rangle\approx 0$.
@@
$\langle\chi_{\mu x}\rangle=\frac{1}{N}\sum_{k}\epsilon(\mathbf{k})\left(\langle f^{\dagger}_{k\mu x\uparrow}f_{k\mu x\uparrow}\rangle+\langle f^{\dagger}_{-k\mu x\downarrow}f_{-k\mu x\downarrow}\rangle\right),$
@@
$\langle\Delta^{\mathbf{x}}_{\mu x}\rangle^{*}=\frac{1}{N}\sum_{k}2\cos{(k_{x})}\langle f^{\dagger}_{k\mu x\uparrow}f^{\dagger}_{-k\mu x\downarrow}\rangle,$